OpenStudy (anonymous):

A curling stone weighing 19.96 kilograms slides down the ice with an initial velocity of 1.7 m/s for 30 seconds. A) how far will the curling stone go before coming to a stop? B) the curling stone hits a second curling stone weighing 18.7 kilograms initially at rest. If the velocity of curling stone A is still 1.7 m/s, what will the final speed and direction of the two curling stones be, assuming it is a perfectly elastic collision.

OpenStudy (roadjester):

I'm actually not sure about part a. For part b, since we're talking about elastic collisions, there is conservation of linear momentum. \[\huge m_1{v_1}_i +m_2{v_2}_i = m_1{v_1}_f +m_2{v_2}_f \]

OpenStudy (anonymous):

really? i thought the equation for elastic collisions was KE\[\frac{ 1 }{ 2 }mv _{Ai}^{2} + \frac{ 1 }{ 2} mv _{bi}^{2} = \frac{ 1 }{ 2 }mv _{Af}^{2} + \frac{ 1 }{ 2} mv _{\bf}^{2}\]

OpenStudy (roadjester):

Oh, forgot to mention that. Ok, yes you are partially correct. When you have a perfectly inelastic collision both energy and linear momentum must be conserved. The equation I wrote was the conservation of linear momentum. What you wrote was the conservation of kinetic energy. You need BOTH to solve this problem. Sorry bout that.

OpenStudy (roadjester):

OOPS i mean elastic

OpenStudy (anonymous):

I'm confused. how do i do this problem then. sorry its just my teacher is really fast and she's expecting us to make up our own problems but we need to solve it and i just don't know how to do this

OpenStudy (roadjester):

ok so you have an ELASTIC collision. when that happens you need to conserve two things: kinetic energy and momentum Momentum: \[\huge m_1{v_1}_i +m_2{v_2}_i =m_1{v_1}_f +m_2{v_2}_f \] Kinetic energy: \[\huge {\frac 1 2} m_1{v_1}_i^2 +{\frac 1 2 } m_2{v_2}_i^2 ={\frac 1 2 } m_1{v_1}_f^2 +{\frac 1 2 } m_2{v_2}_f^2 \] The basic idea is 2 equations 2 unknowns. Make sense?

OpenStudy (anonymous):

sort of. its just that i got weird answers for it. for the final velocity of A i got 0.915 m/s and that doesn't look right

OpenStudy (roadjester):

ok, so are you familiar with the concept needing the same number of equations as you have unknown variables?

OpenStudy (anonymous):

yes

OpenStudy (roadjester):

This is essentially what you need here. Both of the final velocities are unknown so you need two equations to solve them. Conservation of momentum \[\huge m_A{v_A}_i + m_B{v_B}_i =m_A{v_A}_f + m_B{v_B}_f\] Conservation of kinetic energy \[\huge {\frac 1 2}m_A{v_A}_i^2 + {\frac 1 2}m_B{v_B}_i^2 ={\frac 1 2}m_A{v_A}_f^2 + {\frac 1 2}m_B{v_B}_f^2\]

OpenStudy (roadjester):

\[\large 19.96(1.7) + 0 =19.96{v_A}_f + 18.7{v_B}_f\] \[\large {\frac 1 2}(19.96){(1.7)}^2 + 0 ={\frac 1 2}(19.96){v_A}_f^2 + {\frac 1 2}(18.7){v_B}_f^2\]

OpenStudy (roadjester):

Were these your two equations?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

but then it simplified to \[19.96(1.7)^{2} = 19.96 V _{af}^{2}+18.7V_{\bf}^{2}\]

OpenStudy (anonymous):

and then i got \[57.6844= 19.96 vaf ^{2} + 18.7 vbf ^{2}\] which then made 0= -57.6844 +19.96 vaf2 + 18.7 vbf2 then i used the quadratic formula and got vaf= 1.31 m/s

OpenStudy (roadjester):

you can't do that because you're trying to solve two unknowns using one equation. The kinetic energy one. You have to solve both simultaneously

OpenStudy (anonymous):

okay how do u do that

OpenStudy (roadjester):

are you allowed to use a calculator?

OpenStudy (anonymous):

yep

OpenStudy (roadjester):

graphing?

OpenStudy (anonymous):

no

OpenStudy (roadjester):

ok then we'll have to do this the old fashioned way

OpenStudy (anonymous):

oh u mean a graphing calculator yes, sorry. tired

OpenStudy (roadjester):

oh ok then TI- which number do you have?

OpenStudy (anonymous):

84

OpenStudy (anonymous):

84 plus

OpenStudy (roadjester):

aww; oh well; 84 can't solve multi-variable equations only 89 and 92 I think never mind, we can simplify it to a single variable brb

OpenStudy (anonymous):

okay

OpenStudy (roadjester):

\[\large 19.96(1.7) =19.96{v_A}_f + 18.7{v_B}_f => 33.932 = 19.96v_Af+18.7v_Bf\] \[\ {\frac 1 2}(19.96){(1.7)}^2 ={\frac 1 2}(19.96){v_A}_f^2 + {\frac 1 2}(18.7){v_B}_f^2=>28.8422=9.98v_Af^2+9.35v_Bf^2\]

OpenStudy (roadjester):

Take both equations and simplify them as much as possible first.

OpenStudy (roadjester):

Then using the momentum equation, solve for either v_A f or v_B f. Take that and substitute it into the energy equation so that you now have only one variable. From there you can use the solve function on your calculator.

OpenStudy (roadjester):

Make sense?

OpenStudy (roadjester):

by my calculation you should get Va_f = 0.0554m/s Vb_f=1.755 m/s

OpenStudy (anonymous):

i didn't get that

OpenStudy (roadjester):

\[\huge vAf={33.932-18.7vBf\over 19.96}\] did you get that at least? or did you solve for vbf?

OpenStudy (anonymous):

no i got that

OpenStudy (roadjester):

\[\large 28.8422=9.98[{33.932-18.7vBf\over 19.96}]^2 + 9.35vBf^2\]

OpenStudy (roadjester):

this?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

i just don't know where to go from there

OpenStudy (roadjester):

umm, are you any good with algebra?....

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

its just I'm stuck bc i haven't done this in a while

OpenStudy (roadjester):

not sure what to say...it's all just math now

OpenStudy (anonymous):

i got 0.223 bc i took everything in parentheses, squared it, multiplied it all by 9.98. then i made the 9.35 vbf have a common denominator and got 37176.0488vbf2/3976.048 now tell me what i did wrong

OpenStudy (anonymous):

PLEASE. i need this for tomorrow

OpenStudy (roadjester):

one sec; let me finish dinner

OpenStudy (roadjester):

okay let's see

OpenStudy (roadjester):

\[\huge 28.8422 = 9.98({32.932-18.7v_B\over 19.96})^2 +9.35v_B\]

OpenStudy (roadjester):

umm the last term is 9.35v_B

OpenStudy (anonymous):

omg i got it thank u solo much!

OpenStudy (roadjester):

\[\large 28.8422=9.98(0.878v_B^2-3.09v_B+2.72)+9.35v_B\] can you get here at least?

OpenStudy (anonymous):

Part A seems like a dubious question. It might be interpreted as how far will it go in that 30s, otherwise, you'll need a coefficient of friction for stone on ice to find the resisting force.